Prev Next

SW&IT Meeting Minutes 2022-03-14

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20220307
LR20220307 - Bec, Matthieu 40.0
LR20220307 - Mechtley, Brandon 0.0
LR20220307 - Eychaner, Glenn 40.0
LR20220307 - Edmeades, David 0.0
LR20220307 - Grenz, Paul 20.0
LR20220307 - Hooper, Stephen 0.0
LR20220307 - Zhang, Yang 40.0
LR20220307 - Zhang, Xianyu 0.0

News from management

  • Upcoming Changes to COVID-19 Protocol on Face Coverings (email)

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • OC, Management, ATT
    • SHARK-NIR: RFAs
    • AO auxtasks, skipframeTask investigations
    • argos-TwiceAsNice in github
    • safety training
  • This week
    • SHARK-NIR PSR
    • safety training ctd.

YZ

  • Previous:
    • Some MODS and LBC work:
      • A draft MODS image checking program available.
      • Verified all "tmp*" files in /lbt/data/new (total 226 of them, and deleted them after verification).
      • Worked with Olga and checked much of mods2data's historical data (~2800 images).
      • Worked with Olga on studying previously reported LBC DOFPIA problems, and discussions on possible future plans.
    • TMS meeting and some minor software work.
  • Next:
    • Continue improving MODS image checking program (and checking all mods1data historical images).
    • Working with Olga on questions to OSU team related to IT#8567.
    • Update IT#8509 based on recent study.
    • Perhaps reminding Al about the work needed for the archive machine setup.
    • Resume some TMS software work (and possible tests, depending on Heejoo's schedules).

DE

  • Vault VM restore test
    • Looks good, Ketiv agrees.
    • Will remove test VM today
  • Mod backup machine OS to not auto reboot for updates; it already broke one backup
  • Veeam patch for recent RCE; probably not a big risk for us positioned as we are but I made sure to patch it asap regardless
  • New laptop for AC ordered, to be delivered to office
  • DT complaint about space on obs5; lots of logs sitting in /var. I think messages was too big to get rotated, so it just sat there. I changed the config and manually compressed the large messages file so it shouldn't be a problem now
Next:
  • Haven't heard anything from JW re: new laptop, so I'll just install a selection of CAD software and ship it. It will definitely generate some support traffic, but that is inevitable.
  • Nearly automated poking of Veeam re: licensing. We also need to re-up our support for the license we have, but hopefully we can tie all the activity together
  • Resubmit .edu access request to Autodesk (annual thing)
  • Watch for time issues. Both of my phones required a reboot to lock to the right time zone, unsure why.

GE

  • Last week:
    • IT 7493 (Image Subtraction): After deploying LUCI 2.14.8, got bit by the image subtraction problems again. After considerable testing, discovered that image subtraction commands must be issued from the Event Dispatch thread(!). Withdrew 2.14.8 (thanks Stephen!), deployed a new test GUI with several threading and concurrency changes, and reported the bug to Aladin.
    • IT 6818 (LUCI2 Amplifier): D.Thompson pointed out that the correction was off by a row and a row short; fixed. Closed.
    • IT 8326 (Pixel Histogram): Barry noted that the Pixel Histogram failed to appear again. Reported the bug to Aladin.
    • Checked the scripts in lcsp/bin on luci.luci and in stow on observer machines into GitHub.
    • Helped Olga decode a Zemax file so it could be read in numpy.
  • This week
    • Keep testing the new RTD and fix minor bugs as they come up. Hopefully deploy LUCI 2.14.9.
    • Spend a little time looking into alternative image displays, since Aladin seems to be extremely buggy.
    • Restart work on the observer GUI.

PSG

  • Last week
    • Fixed proxy code - ARGOS software was adding extra carriage returns.
    • Fixed Telemetry initialization error in OSS (off by one error)
    • Tested, fixed NOMIC, new errors with initialization.
    • Added features to MACIE code for SHARK-NIR
  • This week
    • NOMIC won't initialize, additional testing needed.
    • Switch back to GCS, get code checked in with new branch.
    • Test new OSS

BM

  • Last week:
    • gain ramp code from Xianyu, AO ops discussion
    • ao arbitrator interruptible commands PR
    • pause/resume gain tests with Juan Carlos, collecting frames
  • This week:
    • finish integrating Xianyu's gain ramp changes
    • ncpa bank tests?
    • interrupted offsets
    • MsgD -RTDB issues:
      • AOS disconnect
      • disturbance loading: AOS dropping notifications of wfsarb.L.MSG
    • Other issues:
      • loopfault during autogain: reset autogain gains to 1
      • Handling BCU timeouts on WFS setup
      • Ice connection lost errors

SH

  • Moving data from old to new Synology
    • LBTI
    • Start on home directories
  • Upgrade switch with LeRoy
  • VMware machine
  • Upgrade 507 computer
  • Tucson network issues

next:

  • VMware installation mountain
  • Mountain network new machines
  • Tucson network issues
Topic revision: r8 - 14 Mar 2022, BrandonMechtley
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback